OpenStudy (anonymous):

What is the volume of 50.0 g of N2 gas at a pressure of 650 mm Hg and a temperature of 15 degrees C?

OpenStudy (stephstar101):

PV=nRT V=(nRT)/P 650mmHg/760mmHg= 0.855atm <-convert mmHg to atmospheres 15+273=288K V=(50.0g/(28.0134g/mol)*0.0821 L-atm/mol-K*288K)/0.855atm V=49.3L